OpenStudy (anonymous):

y= 3x + 3b, how do you solve for b?

OpenStudy (anonymous):

Ok so this means we want to isolate b, I will walk you through the steps: y = 3x + 3b -3x on both sides y -3x = 3b /3 on both sides (y-3x)/3 = b b = (y - 3x)/3 <-- Final answer